OpenStudy (anonymous):

A hot air balloon rises straight up 500 feet, then catches a breeze and is blown west 1,000 feet. How far is the balloon from the point of liftoff

OpenStudy (anonymous):



OpenStudy (anonymous):

I hope this drawing makes sense, let me break it apart. The solid lines are the information you know. Dotted lines are things you need to figure out. You can use the pythagorean theorem to figure out the length of "a." Then from there you'll need to use the pythagorean theorem again to find what (a + 500)² + (a)² = [Your Answer] (this is the dotted line that goes diagonal from the liftoff to the point its currently at.

OpenStudy (mathstudent55):

 d^2 = 500^2 + 1000^2